# HBS - Hardware Build System: A Tcl-based, minimal common abstraction approach for build system for hardware designs

### v3

@@ 5. Hardware Build System Commands
doc Show documentation for hbs Tcl symbol or EDA tool
dump-cores Dump info about cores in JSON format
@@
help Print help message for program or given command
list-cores List cores
list-targets List targets
list-tb List testbench targets
@@
Type hbs help <command>’ to obtain more
information about particular command.
5.1. Viewing HBS API documentation
The doc command was added to ease viewing documentatino for HBS Tcl symbols. Listing 23 presents an example of the doc command usage.
Listing 23: HBS doc command example.
[user@host tmp]\$ hbs doc AddFile
# AddFile adds files to the tool flow. Multiple files with different
# extensions can be added in a single call. args is the list of
# patterns used for globbing files. The file paths are relative to
# the ‘.hbs file path where the proc is called.
proc AddFile {args}
5.2. Dumping cores information
The dump-cores command allows dumping information about found cores into JSON format. The generated JSON data can be used for further processing. For example, the graph command utilizes data from JSON dump to generate the dependency graph.
5.3. Generating dependency graph
The graph command allows generating target dependency graphs in PDF format. Listing 24 shows the graph command help message.
Listing 24: HBS graph command help message.
Output dependency graph for given target.
hbs graph path/to/file/with/cores.json [top-target-path]
If top target path is not provided, then it is assumed to be the
same as the base of the json file path, without the ’.json suffix.
The graph command requires information about cores in the JSON format. This implies that the user must execute the dump-cores or run command before generating a dependency graph. However, this is not a major issue in practice. Dumping cores, even for large designs, does not take more than a few seconds, as the dump-cores command does not run the target tool flow.
Figure 2 presents an example dependency graph generated for VSC8211 (Ethernet PHY) chip tester design.
Figure 2. Example dependency graph generated by HBS graph command.
5.4. Listing cores
The list-cores command allows listing all cores discovered by the hbs program. Listing 25 shows list-cores command help message.
Listing 25: HBS list-cores command help message.
List cores found in .hbs files.
hbs list-cores [core-path-patterns...]
If core-path-patterns are not provided, all found cores are listed.
If core-path-patterns are provided, then only cores whose path
contains at least one pattern are listed.
For example, let us assume the user wants to list all cores in a project related to the APB bus. Listing 26 presents the output of such an action carried out in the VSC8211 tester design.
Listing 26: An example HBS list-cores command output.
[user@host tmp] hbs list-cores apb
vhdl::amba5::apb::bfm
vhdl::amba5::apb::checker
vhdl::amba5::apb::crossbar
vhdl::amba5::apb::mock-completer
vhdl::amba5::apb::pkg
vhdl::amba5::apb::serial-bridge
vhdl::amba5::apb::shared-bus
5.5. Listing targets
The list-targets command allows listing all targets discovered by the hbs program. Listing 27 shows the list-targets command help message. The list-targets command is analogous to the list-cores command but works on targets instead of cores.
Listing 27: HBS list-targets command help message.
List targets for cores found in .hbs files.
hbs list-targets [target-path-patterns...]
If target-path-patterns are not provided, all found targets
are listed. If target-path-patterns are provided, then only
targets whose path contains at least one pattern are listed.
For example, if you want to list all targets for a given core,
simply provide the core path as the argument.
For example, let us assume the user wants to list all targets of the APB serial bridge core. Listing 28 presents the output of such an action carried out in the VSC8211 tester design.
Listing 28: An example HBS list-targets command output.
[user@host tmp] hbs list-targets apb::serial
vhdl::amba5::apb::serial-bridge::src
vhdl::amba5::apb::serial-bridge::tb-write
vhdl::amba5::apb::serial-bridge::tb-read
5.6. Listing testbench targets
The list-tb command allows listing all testbench targets discovered by hbs. The list-tb command is analogous to the list-targets command, but it works on testbench targets instead of all targets. For example, let us assume the user wants to list all testbench targets of the APB serial bridge core. Listing 29 presents the output of such an action carried out in the VSC8211 tester design. As can be seen, the src target is not included.
Listing 29: An example HBS list-tb command output.
[user@host tmp] hbs list-tb apb::serial
vhdl::amba5::apb::serial-bridge::tb-write
vhdl::amba5::apb::serial-bridge::tb-read
5.7. Running target
The run command allows running target procedures. Usually, targets are run to carry out the build process or simulation. However, the user is free to carry out any action in the target being run. Listing 30 shows the run command help message. Running targets is described in subsections 4.5, 4.6, and 4.7.
Listing 30: HBS run command help message.
Run provided target.
hbs run <target-path> [target-args...]
The target path must be absolute target path containing the core
path and target name. Target arguments are forwarded to the target
procedure call.
Example:
hbs run core::path::target-name synthesis
5.8. Running testbench targets
The test command allows running all automatically discovered testbench targets. Listing 31 shows the test command help message.
Listing 31: HBS test command help message.
Run testbench targets.
hbs test [-workers N] [target-path-patterns...]
Testbench targets are detected automatically.
A testbench target is a target which name:
- starts with "tb-" or "tb_",
- ends with "-tb" or "_tb",
- equals "tb".
-workers specifies the number of targets to run simultaneously
(in parallel). N must be a positive integer. -workers must
be provided before target path patterns. Otherwise, it will
be treated as one of the target path patterns. By default,
the number of workers equals the number of available CPUs.
If target path patterns are not provided, all testbench targets
are run. If target path patterns are provided, then only testbench
targets whose path contains at least one pattern are run.
5.9. Locating core definition
The where command allows easily locating .hbs files in which given cores are defined. Listing 32 shows the where command help message.
Listing 32: HBS where command help message.
Print where given cores are defined.
hbs where [core-path-patterns...]
If core-path-patterns are not provided, all found cores are printed.
If core-path-patterns are provided, then only cores whose path
contains at least one pattern are printed.
For example, let us assume the user wants to locate the APB serial bridge core. Listing 33 presents the output of such an action carried out in the VSC8211 tester design.
Listing 33: An example HBS where command output.
[user@host tmp] hbs where apb::serial-bridge
vhdl::amba5::apb::serial-bridge /tmp/vsc8211-tester/gw/apb/apb.hbs
6. Hardware Build System Evaluation
This section tries to evaluate the implemented HBS in terms of other hardware build systems.
Comparing build systems is not straightforward. There are no objective numerical metrics that can be used for build systems comparison. One of such metrics could be the build time. However, in the case of hardware designs, the share of time consumed by EDA tools is orders of magnitude greater than the time overhead introduced by the build system code execution.
Some generic software evaluation metrics could be used, such as numerical or memory complexity. However, the memory overhead introduced by build systems is neglectable compared to the memory utilization of EDA tools. The numerical complexity might be important for the build system developers, but it is rather irrelevant for build system users.
This section tries to evaluate the HBS from some practical viewpoints, such as dependencies, overall internal complexity, ease of use, and provided functionalities.
6.1. Dependencies
The implemented HBS has only the following three dependencies:
  1. (1)
  2. (2)
    Python (optional) - required to run commands not directly related to the build (automatic testbench running and dependency graph generation).
  3. (3)
    Graphviz (optional) - required for the graph command to generate a dependency graph in PDF format.
None of the existing hardware build systems capable of similar features has such a low number of dependencies.
6.2. Overall internal complexity
The measure used to estimate the overall internal complexity of the build system is the number of files, the number of lines of code, and the number of dependencies. The following type of files were excluded from analysis: author files, news files, contributors files, license files, readme files, tests files, and example files.
Table 1 presents metric values for the analyzed hardware build systems. The $\geq N$” signature means at least $N$. For build systems implemented in Python, the number of dependencies was analyzed using the pipreqs (V., 2025) tool. For build systems implemented in Rust, the number of dependencies was extracted from the Cargo.toml files. For flgen and bazel_hdl_rules, the number of dependencies was analyzed manually. Determining the exact number of dependencies (except for HBS) was not straightforward, as there might be additional runtime dependencies. Hence, the need for the greater-than-equal sign.
Besides the complexity metrics, the table also includes a summary of hardware build system features. The Introduced File Extensions is the number of custom file extensions used by a build system. The Introduced File Syntaxes is the number of custom syntaxes used by a build system. For example, FuseSoc scans for files with the .core extension. However, the .core files have a standard YAML syntax. This is why the number of introduced file extensions equals 1, but the number of introduced file syntaxes equals 0.
@@ 6.2. Overall internal complexity
An extra context has to be introduced to the table, as different tools support different features and EDA tools.
@@
Direct Tcl build systems (except HBS) are not included in the table. This is because they support only one EDA tool, and they do not support automatic testbench detection and dependency graph generation. OSVVM-Script is not included because it targets only simulation tools.
@@
The FuseSoc build tool internally utilizes the external Edalize (Kindgren, 2025a) library for interaction with the EDA tools. Including Edalize, the number of files for FuseSoc equals 215, and the number of lines of code equals 19788.
Only HBS, SiliconCompiler, bender, and orbit are capable of generating dependency graphs.
@@
Total number of lines of code of any hardware build system greatly depends on the number of supported EDA tools. FuseSoc supports the greatest number of tools, almost 40. HBS currently supports only five tools (AMD Vivado, AMD xsim, GHDL, GOWIN IDE, nvc). The most complex one is Vivado, which full support requires roughly 220 lines of code. However, supporting the nvc simulator required roughly 170 lines of code. Assuming that HBS will one day support 40 EDA tools, the number of lines of code should still be below 10000.
6.3. Ease of use
@@
The subsection presents the ease of use of HBS in a very common task. Namely, the engineer implements a module and wants to scope some constraints to this module. As an example, a two-stage flip-flop clock domain crossing synchronizer is used. The source code of the synchronizer is not relevant to the analysis. False path is set as a constraint, although the max delay is also frequently used.
@@
Listing 34 presents the synchronizer core definition in FuseSoc.
@@
Listing 34: Cross domain crossing synchronizer core definition in FuseSoc.
@@
Listing 35 presents the content of the vivado-constr.xdc file and Listing 36 presents the content of the constr.tcl file.
Listing 35: Content of the vivado-constr.xdc file.
@@
Listing 36: Content of the constr.tcl file for FuseSoc.
@@
The number of required files is greater than the number of logical elements. The user describes the module in HDL, defines the core for the build system, and adds extra constraints. The number of logical elements equals three, not four. Moreover, the logic of the core definition is more complex to follow, as the user has to switch between files for a very short snippet of code. This leads to a decrease in readability.
Listing 37 presents the same core definition but in the HBS format. The number of required files equals three and is the same as the number of logical elements. The constr.tcl file could be eliminated because the code of the build system is executed directly by the EDA tool. Even the error message is improved. It is possible to report precisely for which EDA tool the constraint file definition is missing. Again, this is possible because the code of the build systems is executed directly by the EDA tool.
In the case of FuseSoc, build scripts are prepared before the EDA tool is run. Within the constr.tcl script, there is no access to the FuseSoc context and variables, as the build system code execution is already finished.
Listing 37: Cross domain crossing synchronizer core definition in HBS.
@@
7. Conclusions
This work proposes a new Tcl-based hardware build system with the following features:
@@ 7. Conclusions
    Low internal code complexity.
@@
    Minimal command line user interface.

### v4

@@ 5.8. Hardware Build System Commands
help Print help message
doc Show documentation for cores
dump Dump info about cores in Tcl dictionary format
dump-json Dump info about cores in JSON format
@@
info Show information on hbs Tcl symbol or EDA tool
ls-cores List cores found in .hbs files
ls-targets List targets for given core
ls-tb List testbench targets
@@
dry-run Run given target without executing and evaluating commands
@@
Type hbs help <command>’ to obtain more information about particular command.
6. Direct-Tcl and indirect-abstract approaches comparison
This section aims to characterize and compare direct-Tcl and indirect-abstract approaches for hardware build systems. The goal of the comparison is not to emerge the best hardware build system, but to present differences between build systems utilizing the direct-Tcl approach (build system code executed directly by EDA tools during the design build flow) and the ones utilizing the indirect-abstract approach (build system generates Tcl script which is later run by a proper EDA tool). The best choice of the hardware build system depends on the project or user requirements.
Comparing hardware build systems is not straightforward. There are no objective numerical metrics that can be used for the build systems comparison. One such metric could be build time. However, in the case of hardware designs, the share of time consumed by EDA tools is orders of magnitude greater than the time overhead introduced by the build system code execution. This is why the build system execution time overhead is not taken into consideration at all.
Some generic software evaluation metrics could be used, such as numerical or memory complexity. However, the memory overhead introduced by build systems is negligible compared to the memory utilization of EDA tools. The numerical complexity might be important for the build system developers, but it is rather irrelevant for build system users.
This section provides a comparative analysis of direct-Tcl and indirect-abstract hardware build systems from a practical perspective. The qualitative assessment examines key operational attributes, including abstraction level, dependency overhead and its potential implications, internal code structure, constraints handling, handling EDA tools without Tcl interface, arbitrary external program calling, extra features, implementation complexity and learning difficulty.
Table 1 presents some numerical and functional features for the analyzed hardware build systems. The $\geq N$” signature means at least $N$. For build systems implemented in Python, the number of dependencies was analyzed using the pipreqs (V., 2025) tool. For build systems implemented in Rust, the number of dependencies was extracted from the Cargo.toml files. For flgen and bazel_hdl_rules, the number of dependencies was analyzed manually. Determining the exact number of dependencies (except for HBS) was not straightforward, as there might be additional runtime dependencies. Hence, the need for the greater-than-equal sign.
The theory of operation of both approaches was analyzed in section 2, and is not included again in this section.
@@ 6. Direct-Tcl and indirect-abstract approaches comparison
Distributed
execution of
compilation flows
N N Y N N N N N N N N
6.1. Abstraction level
The direct-Tcl and indirect-abstract approaches present entirely different paradigms for abstraction over EDA tool interfaces.
In the direct-Tcl approach, the abstraction layer over EDA tool interfaces is limited to the common, primary actions. For example, in HBS, the following actions constitute the common abstraction layer over multiple EDA tools:
  1. (1)
    Target device setting. All synthesis EDA tools require information on the target device. This is why setting the device became part of the common abstraction layer.
  2. (2)
  3. (3)
    Library setting - setting HDL file library.
  4. (4)
    HDL standard setting - setting HDL file standard revision. The build system has to manage this because some tools can not analyze different design units with different standard revisions, for example, nvc (Gasson, 2026). In such a case, the build system must decide which common standard revision to use for analyzing all HDL files.
  5. (5)
  6. (6)
  7. (7)
    Design top module setting - all EDA tools performing simulation or synthesis require information about the top module.
  8. (8)
    Exit severity setting - testbenches are designed to exit with an error code when a specific severity message occurs. This is independent of the simulator being used, hence it should be hidden under the build system abstraction.
In the direct-Tcl approach, the build system intentionally exposes EDA tool-specific quirks so that the user has to resolve them explicitly. This is usually achieved with tool guard clauses. Listing 16 presents an example of such a guard. The snippet presents a helper testbench procedure definition for the AMBA 5 APB crossbar. If the tool is set to the nvc simulator, an extra argument ”–dump-arrays” must be passed to the simulator to dump array signals to the waveform file. This can be seen as a peculiarity of the nvc simulator.
Listing 16: An example of tool guard clause in HBS (direct-Tcl approach).
proc _tb {top} {
hbs::AddDep vhdl::amba5::apb::checker::src
hbs::AddDep vhdl::amba5::apb::bfm::src
hbs::AddDep vhdl::amba5::apb::mock-completer::src
if {\$hbs::Tool eq "nvc"} {
hbs::AddPreSimCb hbs::SetArgSuffix "--dump-arrays"
hbs::SetTop \$top
hbs::AddDep \$hbs::ThisCorePath\::src
In the indirect-abstract approach, the main purpose of the abstraction layer is to provide a great, unified abstraction over EDA tools. As such, the abstraction layer covers many more aspects than the abstraction layer in the direct-Tcl approach. However, maximizing the abstraction layer is not the goal of the indirect-abstract approach; the goal is to unify the user experience.
Instead of exposing EDA tools peculiarities to the user, the indirect-abstract approach tries to hide them under the abstraction layer. For example, the nvc peculiarity of not dumping arrays signals by default, in the indirect-abstract approach is solved in such a way, that build systems automatically add the ”–dump-arrays” argument when simulation is run.
6.2. Dependency overhead and its potential implications
The term ”dependency” in the following context means a program, library, or operating system package that must be explicitly installed by the user to utilize all features of the analyzed build system. Any tools installed automatically alongside EDA tools are not treated as dependencies.
Hardware build systems representing the direct-Tcl approach tend to have fewer dependencies. For example, the implemented HBS has only the following two dependencies:
  1. (1)
  2. (2)
    Graphviz (optional) - required for generating a core dependency graph in a graphical format.
None of the existing hardware build systems representing the indirect-abstract approach, capable of similar features, has so few dependencies.
The explanation for the lower number of dependencies in build systems that use the direct-Tcl approach is straightforward. They rely on libraries and programs that are installed during EDA tool installation. On the other hand, build systems utilizing the indirect-abstract approach often rely on dependencies installed manually via package managers.
Nevertheless, a pure number of dependencies might not be a good metric for comparison. A project could have a single dependency that suffers from multiple issues, and another project could have hundreds of well-maintained dependencies.
What might be more important from the user’s point of view is stability. A build system is expected to work across many systems with a wide range of Long Term Support. The ease of initial setup of the build system might also be relevant to the user.
The direct-Tcl approach with fewer dependencies is probably less susceptible to stability issues. However, overall stability depends on multiple factors, such as server stability, OS stability, network reliability, CI/CD complexity, etc.
6.3. Internal code structure
The internal code structure was analyzed in terms of the number of files, lines of code, introduced file syntaxes and extensions, and supported features. The following types of files were excluded from analysis: author files, news files, contributors files, license files, readme files, tests files, and example files. The goal of the analysis is to provide raw facts.
In Table 1, the introduced file extensions are the number of custom file extensions used by a build system. The introduced file syntaxes are the number of custom syntaxes used by a build system. For example, FuseSoc scans for files with the .core extension. However, the .core files have a standard YAML syntax. This is why the number of introduced file extensions equals 1, but the number of introduced file syntaxes equals 0.
@@ 6.3. Internal code structure
An extra context has to be introduced to the table, as different tools support different features and EDA tools. Direct-Tcl build systems (except HBS) are not included in the table. This is because they support only one EDA tool, and they do not support automatic testbench detection and dependency graph generation. OSVVM-Script is not included because it targets only simulation tools.
@@
The FuseSoc build tool internally uses the external Edalize (Kindgren, 2025a) library to interact with EDA tools. Including Edalize, the number of files for FuseSoc equals 215, and the number of lines of code equals 19788.
@@
The total number of lines of code of any hardware build system depends heavily on the number of supported EDA tools. FuseSoc supports the greatest number of tools, almost 40. HBS currently supports the following seven tools: AMD Vivado, AMD xsim, GHDL (Gingold, 2026), GOWIN IDE, Intel/Altera Quartus, nvc, Questa. The most complex one is Vivado, which full support requires roughly 190 lines of code. However, supporting the nvc simulator required roughly 170 lines of code. Assuming that HBS will one day support 40 EDA tools, the number of lines of code should remain below 10000.
Tools following the direct-Tcl approach tend to have more monolithic code structure, fewer files, and fewer lines of code. The code for handling EDA tools is often placed in a single file. The lower number of lines of code results from the direct access to the EDA tool commands. There is no need for wrapper code. They prefer shared global variables which makes build logic easier to follow but also more bug-prone.
Tools following the indirect-abstract approach tend to have a more modular code structure. They have more files and a greater number of lines of code. The code for handling specific EDA tools is organized into modules in separate files. The encapsulation makes the code more bug-resistant, but the build logic is slightly harder to follow.
6.4. Constraints handling
Building any hardware design (FPGA or ASIC) without constraints is impossible. Constrains usually define clocks, path delays, timing relations, and physical properties, for example, pin locations, drive strength or voltage level. Any hardware build system must support various constraint types and scenarios. The goal of this subsection is to present differences in constraint handling from the user perspective of the build system.
In the case of a simple constraints scenario with global constraints, there is no difference between the direct-Tcl and indirect-abstract approaches. The user simply defines constraints in a supported file format (e.g., .sdc, .xdc, or .tcl) and adds the file to the project or build flow. Based on the file extension, build systems decide how to handle the file. However, there is a noticeable difference when a project requires constraints scoped to modules and cells, or when constraints are added optionally.
Scoping constraints to module
@@ 6.4. Constraints handling
Let us analyze a simple example. The engineer implements a module and wants to scope some constraints to this module. As an example, a two-stage flip-flop clock domain crossing synchronizer is used. The source code of the synchronizer is not relevant to the analysis. The false path is set as a constraint, although the max delay is also frequently used.
@@
Listing 17 presents the synchronizer core definition in FuseSoc, which represents the declarative indirect-abstract subcategory.
@@
Listing 17: Cross domain crossing synchronizer core definition in FuseSoc.
@@
Listing 18 presents the content of the vivado-constr.xdc file, and Listing 19 presents the content of the constr.tcl file.
Listing 18: Content of the vivado-constr.xdc file.
@@
Listing 19: Content of the constr.tcl file for FuseSoc.
@@
The number of required files (4) is greater than the number of logical elements (3). The user describes the module in HDL, defines the core for the build system, and adds extra constraints. The number of logical elements equals three, and the number of files equals four. The logic of the core definition is more complex to follow, as the user has to switch between files for a very short snippet of code. This may lead to a decrease in productivity (Meyer et al., 2014).
Listing 20 presents the same core definition but in the direct-Tcl approach. The number of required files is three and equals the number of logical elements. The constr.tcl file could be eliminated because the code of the build system is executed directly by the EDA tool. Even the error message is improved. It is possible to report which EDA tool the constraint file definition is missing for. Again, this is possible because the build system code is executed directly by the EDA tool.
In the case of the indirect-abstract approach, build scripts are prepared before the EDA tool is run. Within the constr.tcl script, there is no access to the build system context or variables, as the build system code execution is already finished.
Listing 20: Cross domain crossing synchronizer core definition in HBS (direct-Tcl approach).
@@
In the case of the programmable indirect-abstract subcategory, the solution depends on the tool. For example, in hdlmake, the intermediate Tcl script is also required to scope a constraint file to a module. The same script is also required for the SiliconCompier. Hog, for example, instead of a Tcl script, requires a custom constraint list file (a file with the .con extension) with paths to constraint files and their property declarations. Moreover, the only way to opt out of the default constraints is to comment out the constraint lines in the constraint list file. To the author’s knowledge, none of the existing indirect-abstract build systems utilizes the concept of direct injection of Tcl code into the build script. They all need one additional file (a Tcl script or a tool-specific file) for scoping constraints.
Scoping constraints to cells and optional constraints
Let us analyze a more complex example. Namely, the Advanced Peripheral Bus Clock Domain Crossing bridge (APB CDC bridge). Such a bridge requires some timing constraints due to the clock domain crossing. There are a few possible ways to define valid constraints for the APB CDC bridge. The core designer may want to provide some default constraints. However, the user should also be able to provide their own constraints.
One possible constraint strategy is to set a false path for handshake signals and a maximum delay for bus signals crossing domains. The max delay value should equal the period of the destination clock. The false path constraints can be safely scoped to the APB CDC bridge module. However, because there may be multiple APB CDC bridges in the design crossing between various clock domains, the max delay constraints should be scoped to cells. An alternative solution is to pick the shortest clock period as the value for the max delay constraint for all bridges. However, such an approach leads to design overconstraint and imposes unnecessary requirements on the router.
Listing 21 shows how APB CDC bridge constraints are handled in the HBS, which represents the direct-Tcl approach. The code is concise and located within a single procedure. The user can pass the -no-constr flag to the target to omit adding the default constraints. The flag is checked with a simple if statement. Moreover, if an invalid argument is passed, an error with a precise message is returned. The apb-cdc-bridge-ref.xdc and apb-cdc-bridge-cell.tcl constraint files are not presented as their content is irrelevant. The first one has constraints scoped to the module; the second one has code that applies constraints scoped to cells.
Listing 21: Advanced Peripheral Bus Clock Domain Crossing bridge target definition in HBS (direct-Tcl).
# Use -no-constr flag to omit adding the default constraints.
proc src {args} {
hbs::AddDep vhdl::amba5::apb::pkg::src
hbs::SetLib "lapb"
hbs::AddFile cdc-bridge.vhd
set argCount [llength \$args]
if {\$argCount == 1} {
set arg [lindex \$args 0]
if {\$arg eq "-no-constr"} {
return
} else {
error "\$hbs::ThisTargetPath: invalid arg ’\$arg’, valid args are: ’-no-constr’"
} elseif {\$argCount > 1} {
error "\$hbs::ThisTargetPath: invalid arg count \$argCount, expected at most 1"
if {\$hbs::Tool eq "vivado-prj"} {
hbs::AddFile vivado/apb-cdc-bridge-ref.xdc
set_property SCOPED_TO_REF APB_CDC_Bridge [get_files apb-cdc-bridge-ref.xdc]
hbs::AddFile vivado/apb-cdc-bridge-cell.tcl
set_property USED_IN_SYNTHESIS false [get_files apb-cdc-bridge-cell.tcl]
} else {
error "\$hbs::ThisCore misses constraints for \$hbs::Tool"
How applying the same constraints in the indirect-abstract approach looks depends on the tool subcategory. For example, in FuseSoc, which represents the declarative subcategory, applying the module and cell-scoped constraints would be similar to what was presented in listings 17 and 19 for the CDC synchronizer. Again, one more Tcl file would be required. The optionality of the constraints would be handled using the flags mechanism (Kindgren, 2025b). While the flag mechanism allows conditional inclusion of files, it does not allow arbitrary actions to be executed due to its declarative nature.
In hdlmake, the same core description would require two Manifest.py files. The first one with constraints included, and the second one without. The user would pick a proper Manifest.py file when declaring dependency on the core. With just a single boolean parameter, this approach does not seem complex. However, with multiple parameters, the number of Manifest.py files would equal the number of possible combinations.
In the programmable subcategory of the indirect-abstract approach, the complexity of constraint handling is lower. In this subcategory, the solution is very similar to the one present in the direct-Tcl approach. However, instead of directly injecting the Tcl code into the generated Tcl build script, the user must place it in a separate Tcl file. There is only minor overhead resulting from:
@@
    build system programming language Application Programming Interface (API) call,
@@
    build logic fragmentation due to the requirement of putting the Tcl code in a separate file.
6.5. Handling EDA tools without Tcl interface
Not all EDA tools have a Tcl interface. For example, free and open-source simulators, for example, Verilator (Snyder, 2026) or GHDL, have a shell interface. The user simply calls programs within the shell providing appropriate arguments.
Handling EDA tools without the Tcl interface in the direct-Tcl approach is straightforward. The Tcl language makes it easy to call any external program using the exec command. Listing 22 presents how the elaboration stage for the GHDL simulator is implemented in HBS. The command to be executed is constructed directly as a string to be passed to the shell. The code related to the working directory change and error checking takes more lines than the code creating the actual command. The hbs::Exec procedure is a simple wrapper around the Tcl exec command, which allows printing the command to the standard output in the dry-run mode.
Listing 22: GHDL elaboration stage implementation in HBS (direct-Tcl approach).
proc elaborate {} {
set workDir [pwd]
cd \$hbs::RunTargetBuildDir
set cmd "ghdl -e \$hbs::ArgPrefix --std=[hbs::ghdl::std] --workdir=work \
\$hbs::ghdl::libs \$hbs::ArgSuffix \$hbs::Top"
set err [hbs::Exec \$cmd]
if {\$err} {
hbs::panic "\$hbs::Top elaboration failed with exit status \$err"
cd \$workDir
In the indirect-abstract approach, handling EDA tools without the Tcl interface is implemented differently. In the indirect-abstract approach, the build system code is run before the build code. This implies that the build system code must generate the build code that is only later run. The build code is usually generated in the form of a Makefile or shell script. Compared to the direct-Tcl approach, there are two essential differences: